A 130.0 −mL buffer solution is 0.110 M in NH3 and 0.135 M in NH4Br. The Kb value for NH3 is 1.76×10−5.

If the same volume of the buffer were 0.270 M in NH3 and 0.390 M in NH4Br, what mass of HCl could be handled before the pH falls below 9.00?
Answer #1

So we have a buffer with 130.00mL and with

NH3 + HCl --------------> NH4Br

130.0x0.27=35.1 --- 130.0x 0.390=50.7 initial mmoles

------------ p ------- change

35.1-p 0 50.7+p equilibrium

The pH of this new buffer = 9.00 or less.

Hence pOH of buffer = 14-9.00 = 5.00

pKb of NH3 = -log 1.76x10-5 = 4.7544

thus pH of buffer by Hendersen equation is

pOH = pKb + log [conjugate acid]/[base]

5.00 = 4.7544 + log [50.7+p]/[35.1-p]

solving p = 4.01 mmol

mass of HCl = molx molar mass = 4.01x10-3 mol x36.5g/mol

= 0.146g of HCl